# Transductive Sample Complexities Are Compact

### v2

@@ 1.1 Contributions
Our results involve comparing a hypothesis class $\mathcal{H}$ to its “finite projections.” Formally, for a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and finite collection of unlabeled data $S\subseteq\mathcal{X}$, we refer to the finite subsets of $\mathcal{H}|_{S}$ as finite projections of $\mathcal{H}$. Note that $\mathcal{H}$ is being “made finite” at two levels: first by restricting its functions to a finite region $S\subseteq\mathcal{X}$ of the domain, and second by passing to a finite subset of $\mathcal{H}|_{S}$.
@@
Our results hold for improper learners, i.e., learners that are permitted to emit a predictor outside the underlying class $\mathcal{H}$. Curiously, compactness of sample complexity can be seen to fail strongly when one requires that learners be proper, using the work of Ben-David et al. (2019). This demonstrates a structural difference between proper and improper learning; see Section 3.5 for further detail.
@@ 1.2 Related Work
Perhaps most related to the present work is Attias et al. (2023), which introduces the $\gamma$-OIG dimension and demonstrates that it characterizes learnability for supervised learning problems with pseudometric losses. Notably, this is the first general dimension characterizing learnability across essentially the entirety of supervised learning. The $\gamma$-OIG dimension itself establishes a qualitative form of compactness — as it is defined using only the finite projections of a class — but we note that it has not been shown to tightly characterize the sample complexity of learning. Furthermore, it is analyzed only for realizable learning, which is in general not equivalent to agnostic learning (e.g., for regression). Our work, in contrast, establishes exact compactness for the sample complexity of transductive learning for both the realizable and agnostic settings, with respect to a general class of loss functions. Moreover, in Section 3.5 we extend our results to certain cases of distribution-family learning, including realizable learning of partial concept classes.
@@ 2.2 Transductive Learning
The transductive error rate of a learner $A$ with respect to a class $\mathcal{H}$ is the function $\xi_{A,\mathcal{H}}(n)=\sup_{S\in\mathcal{X}^{n},\;h\in\mathcal{H}}L_{S,h}^{\operatorname{Trans}}(A)$. The transductive sample complexity of $A$ is the function
$m_{\operatorname{Trans},A}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{A,\mathcal{H}}(m^{\prime})\leq\epsilon,\;\forall m^{\prime}\geq m\}.$
We now define the transductive error rate and sample complexity of a hypothesis class as the best-case over all its learners.
@@ 3 Compactness of Learning
We present the central result of the paper in this section: the transductive sample complexity of learning is a compact property of a hypothesis class. In Section 3.1 we study compactness of realizable supervised learning over proper loss functions, and demonstrate a strong compactness result: a class $\mathcal{H}$ is learnable with transductive sample complexity $m$ if and only if all its finite projections are learnable with the same complexity. In Section 3.2 we examine the case of realizable supervised learning over improper loss functions and prove a negative result: the previous compactness result no longer holds in this more general setting. Nevertheless, we demonstrate an approximate form of compactness, up to a factor of 2, for (improper) metric losses. Moreover, we show exact compactness for the special case of the (improper) 0-1 loss function, i.e., multiclass classification over arbitrary, possibly infinite label sets. Notably, this recovers M. Hall’s classic matching theorem for infinite graphs (Hall Jr, 1948) as a corollary to our central result. In Section 3.3 we examine analogues of our results for agnostic learning, and in Section 3.4 we transfer our results to the PAC model via standard equivalences, obtaining approximate compactness of sample complexities. Finally, in Section 3.5 we extend our results to certain forms of distribution-family learning.
@@ 3.1 Realizable Learning With Proper Loss Functions
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. Before arguing the reverse direction, some terminology: a partial assignment of variables is an assignment of variables for a subset of $L$. A partial assignment is said to be completable with respect to $R^{\prime}\subseteq R$ if its unassigned variables can all be assigned so that the functions $r^{\prime}\in R^{\prime}$ are kept below $\epsilon$. A partial assignment is finitely completable if it is completable with respect to all finite subsets of $R$. Notably, this is a pointwise condition, not a uniform one; the completions are permitted to vary across $R$’s subsets.
@@
Thus $f_{R^{\prime}}^{-1}([0,\epsilon]^{k})$ is compact, as is its projection onto its first coordinate. That set is precisely $N_{R^{\prime}}(\ell)$, demonstrating our intermediate claim. We thus have a family of compact, non-empty sets =$\mathcal{I}=\left\{N_{R^{\prime}}(\ell):R^{\prime}\subseteq R,|R^{\prime}|\leq\infty\right\}$. Note that finite intersections of elements of $\mathcal{I}$ are non-empty, as $\bigcap_{i=1}^{j}N_{R_{i}}(\ell)\supseteq N_{\bigcup_{i=1}^{j}R_{i}}(\ell)\neq\emptyset$.
@@
We will capture the task of transductively learning $\mathcal{H}$ on such instances by way of a certain collection $R$ of functions and $L$ of variables. Each variable $\ell\in L$ will be permitted to take values in $\mathcal{Y}$, while each function $r\in R$ depends upon exactly $n$ variables in $L$ and outputs values in $\mathbb{R}_{\geq 0}$. More precisely, let $R=\mathcal{H}|_{S}$ and $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$. These serve merely as representations for the functions in $R$ and variables in $L$, not their true definitions (which will be established shortly). Note now that by suppressing the unlabeled datapoints of $S$, we can equivalently represent elements of $R$ as sequences in $\mathcal{Y}^{n}$ and elements of $L$ as sequences in $(\mathcal{Y}\cup\{?\})^{n}$. In this view, each element of $L$ is precisely an element of $R$ which had exactly one entry replaced with a ``?"``?"``?"` ` ? ".
@@ 3.2 Realizable Learning With Improper Loss Functions
Let us describe the main idea of Theorem 3.8, whose proof is deferred to Appendix B.1. The crucial step lies in the creation of the label space $\mathcal{Y}=R\cup S$, where $R$ is an infinite set whose points are all distance 2 apart, and $S$ is an infinite set whose elements are indexed by the finite subsets of $R$, e.g., as in $s_{R^{\prime}}\in S$ for finite $R^{\prime}\subseteq R$. For all such $R^{\prime}$, define $s_{R^{\prime}}$ to be distance 1 from the elements of $R^{\prime}$, distance 2 from the other elements of $R$, and distance 1 from all other points in $S$. Then $\mathcal{Y}$ indeed forms a metric space, and it is straightforward to see that, for instance, the class of all functions from a one-element set to $\mathcal{Y}$ is more difficult to learn than its finite projections (equivalently, finite subsets).
@@
The proof of Theorem 3.9 is deferred to Appendix B.2, but let us briefly sketch the main idea. Fix $n\in\mathbb{N}$, $S\in\mathcal{X}^{n}$, and set $\epsilon=\xi(n)$. Consider again the collection of functions $R=\mathcal{H}|_{S}$ and variables $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$, as described in the proof of Theorem 3.6. By the premise of the theorem, for any finite subset $R^{\prime}\subseteq R$, there exists an assignment of variables $L\to\mathcal{Y}$ which maintains all functions in $R^{\prime}$ below $\epsilon$. Each such assignment induces an apportionment of error to each function in $R^{\prime}$, i.e., a vector of length $n$ with positive entries summing to $\epsilon$. For $r\in R^{\prime}$ depending upon variables $\ell_{1},\ldots,\ell_{n}$, this apportionment tracks the contribution of each $\ell_{i}$ to the output of $r$. The central technical step of the proof is to demonstrate that one can assign apportionments to each node $r\in R$ such that any finite subset of the apportionments can be satisfied by an assignment of variables $L\to\mathcal{Y}$. Then let $\ell=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ be a variable. We assign $\ell$ to the value $\hat{y}$ such that the function $(y_{1},\ldots,y_{i},\hat{y},y_{i+1},\ldots,y_{n})$ has minimal budget apportioned to $\ell$, among all such $\hat{y}$. From an invocation of the triangle inequality, this learner at most doubles the output of any $r\in R$.
@@
M. Hall’s theorem states that an infinite bipartite graph $G$ is $R$-matchable if and only if it is finitely $R$-matchable, provided that all nodes in $R$ have finite degree. Before proving M. Hall’s theorem by way of Theorem 3.3, we establish an intermediate lemma: informally, all nodes in $G$ can be taken to have finite degree, not merely those in $R$.
@@
The proof of Lemma 3.12 is deferred to Appendix B.3, but its intuition is fairly simple: by P. Hall’s theorem, $G$ is finitely $R$-matchable precisely when Hall’s condition holds, i.e., $|N(R^{\prime})|\geq|R^{\prime}|$ for all finite $R^{\prime}\subseteq R$ (Hall, 1935). Thus any $\ell\in L$ which is not incident to a Hall blocking set can be removed from $G$ while preserving Hall’s condition and finite $R$-matchability. Proceeding in this way, nodes can be removed until each remaining $\ell\in L$ is contained in a Hall blocking set $R^{\prime}_{\ell}$. At this point, $\ell$’s incident edges can be safely restricted to those which are incident with $R^{\prime}_{\ell}$, a finite set.
@@ 3.3 Agnostic Learning
Furthermore, one can use nearly identical reasoning as in the proofs of Theorems 3.6 and 3.7 to see that agnostic transductive learning is described by a system of variables $L$ and functions $R$. In particular, set $R=\mathcal{Y}^{n}$ and let $L\subseteq(\mathcal{Y}\cup\{?\})^{n}$ contain all sequences with exactly one ``?"``?"``?"` ` ? ". Then a function $r=(y_{1},\ldots,y_{n})$ depends upon the variables $\{\ell_{i}\}_{i\in[n]}$, where $\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ and
$r(\ell_{1},\ldots,\ell_{n})=\frac{1}{n}\sum_{i\in[n]}d_{\mathcal{Y}}(y_{i},\ell_{i})-\inf_{h\in\mathcal{H}}\frac{1}{n}\sum_{i\in[n]}\ell(h(x_{i}),y_{i}).$
@@ 3.4 PAC Learning
Proof.
Let $m_{\operatorname{Exp},\mathcal{H}}$ denote the sample complexity of learning $\mathcal{H}$ in the expected error regime (i.e., $m_{\operatorname{Exp},\mathcal{H}}(\epsilon)$ equals the number of datapoints needed to incur expected error at most $\epsilon$). Then we have
The first inequality follows immediately from the fact that the loss function is bounded above by 1. The second inequality follows from a repetition argument, i.e., a learner attaining expected error $\leq\epsilon/2$ on samples of size $n$ can be boosted to attain expected error $\leq\epsilon$ with probability $\geq 1-\delta$ by using an additional factor of $O(\log(1/\delta))$ many samples, as described in (Daniely and Shalev-Shwartz, 2014).
The first inequality follows inequality from a standard leave-one-out argument of Haussler et al. (1994). The second inequality follows from the fact that for any $S=((x_{1},y_{1}),\ldots,(x_{n},y_{n}))$, there exists an $m_{n}\in\mathbb{N}$ such that $m_{n}$ many independent draws from the uniform distribution over the entries of $S$ has probability at least $\frac{1}{\epsilon}$ of containing exactly $n-1$ elements of $S$, as detailed in (Asilis et al., 2024, Lemma A.1). The claim follows from both of the established chains of inequalities. ∎
@@
Let $\mathcal{X}$ be a domain, $\mathcal{Y}$ a label set, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Suppose that the loss function $\ell$ is bounded to $[0,1]$ and satisfies either of the following conditions:
@@
3.5 Distribution-family Learning
The analysis of PAC learning with respect to more flexible distribution classes than the realizable and agnostic cases falls largely under the purview of distribution-family learning (Benedek and Itai, 1991). Formally, a problem in distribution-family learning of a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ is defined by a family of distributions $\mathbb{D}$ over $\mathcal{X}$, such that unlabeled datapoints are drawn from a distribution $D\in\mathbb{D}$ and labeled by a hypothesis $h\in\mathcal{H}$ (in the realizable case) or arbitrarily (in the agnostic case).
Notably, distribution-family learning has infamously resisted any characterization of learnability, combinatorial or otherwise, for the 40 years since its inception. In fact, there is some evidence to suggest that no such characterization may exist (Lechner and Ben-David, 2023). Furthermore, it is a setting in which uniform convergence fails to characterize learning, rendering ineffective many of the standard and most celebrated techniques of learning theory.
Definition 3.21.
Definition 3.21 is sufficiently flexible that we may apply it to distribution-family learning with $\mathcal{Z}=\mathcal{X}$ or to PAC learning over arbitrary distribution classes $\mathbb{D}$ with $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. Though it may appear overly restrictive at first glance, note that well-behavedness is satisfied not only for ordinary PAC learning in the agnostic and realizable cases, but also for learning of partial concept classes in the realizable case (Alon et al., 2022) and for the EMX learning of Ben-David et al. (2019). In particular, though EMX learning is not presented as a supervised learning problem in Ben-David et al. (2019), it can be seen as a binary classification problem over a domain $\mathcal{X}$ for which $\mathcal{H}$ consists of those functions outputting finitely many 1’s and $\mathbb{D}$ contains all realizable, discrete distributions placing all $\mathcal{Y}$-mass on the label 1.
Proposition 3.22.
Proof.
The argument follows identically as in Lemma 3.19. Crucially, $m_{\operatorname{Trans},\mathcal{H}}(\epsilon)\leq m_{\operatorname{Exp},\mathcal{H}}(\epsilon/e)$ due to the fact that a learner attaining low expected error must perform well on the uniform distribution over any particular learning instance $S=(x_{i},y_{i})_{i\in[n]}$ — owing to well-behavedness of $\mathbb{D}$ — and can thus be used to create a good transductive learner, as described in Lemma 3.19. ∎
Note that Proposition 3.22 holds for the natural definition of transductive learning with respect to $\mathbb{D}$, i.e., in which the adversary must select a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$ which lie in the support of some $D\in\mathbb{D}$. It is now immediate from the proof of Theorem 3.6 that distribution-family learning is a setting in which the transductive sample complexity of learning is $\mathcal{H}$ equals the sample complexity of learning its most challenging finite projections. In the following, we let $\mathbb{D}|_{X}\subseteq\mathbb{D}$ denote the distributions of $\mathbb{D}$ which place full measure on $X\subseteq\mathcal{X}$.
Theorem 3.23.
  • 1.
  • 2.
As in Section 3.4, Proposition 3.22 applied to Theorem 3.23 immediately yields an almost-exact form of compactness for distribution-family PAC learning with realizable and well-behaved distribution classes. This demonstrates that the learnability of even problems as exotic as EMX learning can be detected by examining all their finite projections, provided that no restrictions are placed upon learners. In Ben-David et al. (2019), however, learners were required to only emit hypotheses in $\mathcal{H}$. Our work demonstrates that the nature of their undecidability result — in which the learnability of a class $\mathcal{H}$ is determined entirely by its cardinality, despite all its projections being easily learned — could otherwise not appear in supervised learning with metric losses.

### v3

# Transductive Learning Is Compact

@@ 1.1 Contributions
Our results involve comparing a hypothesis class $\mathcal{H}$ to its “finite projections.” Formally, for a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ and any finite collection of unlabeled data $S\subseteq\mathcal{X}$, we refer to the finite subsets of $\mathcal{H}|_{S}$ as finite projections of $\mathcal{H}$. Note that $\mathcal{H}$ is being “made finite” at two levels: first by restricting its functions to a finite region $S\subseteq\mathcal{X}$ of the domain, and second by passing to a finite subset of $\mathcal{H}|_{S}$. Thus, any finite projection of $\mathcal{H}$, e.g. $\mathcal{F}\subseteq\mathcal{H}|_{S}$, is necessarily a finite set of behaviors, $|\mathcal{F}|<\infty$, regardless of whether $\mathcal{H}|_{S}$ in its totality is infinite (as may easily occur if $\mathcal{Y}$ is infinite).
@@
Our results hold for improper learners, i.e., learners that are permitted to emit a predictor outside the underlying class $\mathcal{H}$. Curiously, compactness of sample complexity can be seen to fail strongly when one requires that learners be proper, using the work of Ben-David et al. (2019). This demonstrates a structural difference between proper and improper learning; see Appendix B for further detail.
@@ 1.2 Related Work
Perhaps most related to the present work is Attias et al. (2023), which introduces the $\gamma$-OIG dimension and demonstrates that it characterizes learnability for supervised learning problems with pseudometric losses. Notably, this is the first general dimension characterizing learnability across essentially the entirety of supervised learning. The $\gamma$-OIG dimension itself establishes a qualitative form of compactness — as it is defined using only the finite projections of a class — but we note that it has not been shown to tightly characterize the sample complexity of learning. Furthermore, it is analyzed only for realizable learning, which is in general not equivalent to agnostic learning (e.g., for regression). Our work, in contrast, establishes exact compactness for the sample complexity of transductive learning for both the realizable and agnostic settings, with respect to a general class of loss functions. Moreover, in Appendix B we extend our results to certain cases of distribution-family learning, including realizable learning of partial concept classes.
@@ 2.2 Transductive Learning
The transductive error rate of a learner $A$ for $\mathcal{H}$ is the function $\xi_{A,\mathcal{H}}:\mathbb{N}\to\mathbb{R}$ defined by $\xi_{A,\mathcal{H}}(n)=\sup_{S\in\mathcal{X}^{n},\;h\in\mathcal{H}}L_{S,h}^{\operatorname{Trans}}(A)$. The transductive sample complexity of a learner $A$ for $\mathcal{H}$ is the function $m_{\operatorname{Trans},A}(\epsilon)=\min\{m\in\mathbb{N}\ :\ \xi_{A,\mathcal{H}}(m^{\prime})\leq\epsilon,\;\forall m^{\prime}\geq m\}$.
@@
We say that $\mathcal{H}$ is learnable in the realizable case with transductive sample function $m$ when $m_{\operatorname{Trans},\mathcal{H}}(\epsilon)\leq m(\epsilon)$ for all $\epsilon$.
@@ 3 Compactness of Learning
We present the central result of the paper in this section: the transductive sample complexity of learning is a compact property of a hypothesis class. In Section 3.1 we study compactness of realizable supervised learning over proper loss functions, and demonstrate a strong compactness result: a class $\mathcal{H}$ is learnable with transductive sample complexity $m$ if and only if all its finite projections are learnable with the same complexity. In Section 3.2 we examine the case of realizable supervised learning over improper loss functions and prove a negative result: the previous compactness result no longer holds in this more general setting. Nevertheless, we demonstrate an approximate form of compactness, up to a factor of 2, for (improper) metric losses. Moreover, we show exact compactness for the special case of the (improper) 0-1 loss function, i.e., multiclass classification over arbitrary, possibly infinite label sets. Notably, this recovers M. Hall’s classic matching theorem for infinite graphs (Hall Jr, 1948) as a corollary to our central result. In Section 3.3 we examine analogues of our results for agnostic learning, and in Section 3.4 we transfer our results to the PAC model via standard equivalences, obtaining approximate compactness of sample complexities. Due to space constraints, we defer an extension of our results to distribution-family PAC learning to Appendix B.
@@ 3.1 Realizable Learning With Proper Loss Functions
(1.)(2.)(1.$\implies$(2.)( 1 . ) ⟹ ( 2 . ) is immediate. Before arguing the reverse direction, some terminology: a partial assignment of variables is an assignment of variables for a subset of $L$. A partial assignment is said to be completable with respect to $R^{\prime}\subseteq R$ if its unassigned variables can all be assigned so that all functions $r^{\prime}\in R^{\prime}$ are kept below $\epsilon$. A partial assignment is finitely completable if it is completable with respect to all finite subsets of $R$. This is a pointwise condition: the completions are permitted to vary across $R$’s subsets.
@@
Thus $f_{R^{\prime}}^{-1}([0,\epsilon]^{k})$ is compact, as is its projection onto its first coordinate. That set is precisely $N_{R^{\prime}}(\ell)$, demonstrating our intermediate claim. We thus have a family of compact, non-empty sets =$\mathcal{I}=\left\{N_{R^{\prime}}(\ell):R^{\prime}\subseteq R,|R^{\prime}|<\infty\right\}$. Note that finite intersections of elements of $\mathcal{I}$ are non-empty, as $\bigcap_{i=1}^{j}N_{R_{i}}(\ell)\supseteq N_{\bigcup_{i=1}^{j}R_{i}}(\ell)\neq\emptyset$.
@@
$(0,0,?)$
$L$
$(1,0,?)$$(0,1,?)$$(0,?,0)$$(1,?,0)$(?,0,0)?00(?,0,0)( ? , 0 , 0 )(?,1,0)?10(?,1,0)( ? , 1 , 0 )(0,0,0)000(0,0,0)( 0 , 0 , 0 )
$R$
$(1,0,0)$$(0,1,0)$
Figure 1: Depiction of variables $L$ and functions $R$ which model transductive learning, for a sequence of unlabeled datapoints $|S|=3$ such that $\mathcal{H}|_{S}$ contains the behaviors $(0,0,0)$, $(1,0,0)$, and $(0,1,0)$. Arrows denote functional dependence, i.e., each $r\in R$ depends upon its incident variables.
@@
We will capture the task of transductively learning $\mathcal{H}$ on such instances by way of a certain collection $R$ of functions and $L$ of variables. Each variable $\ell\in L$ will be permitted to take values in $\mathcal{Y}$, while each function $r\in R$ depends upon exactly $n$ variables in $L$ and outputs values in $\mathbb{R}_{\geq 0}$. More precisely, let $R=\mathcal{H}|_{S}$ and $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$. These serve merely as representations for the functions in $R$ and variables in $L$, not their true definitions (which will be established shortly). Note now that by suppressing the unlabeled datapoints of $S$, we can equivalently represent elements of $R$ as sequences in $\mathcal{Y}^{n}$ and elements of $L$ as sequences in $(\mathcal{Y}\cup\{?\})^{n}$. In this view, each element of $L$ is precisely an element of $R$ which had exactly one entry replaced with a ``?"``?"``?"` ` ? ". See Figure 1.
@@ 3.2 Realizable Learning With Improper Loss Functions
Let us describe the main idea of Theorem 3.8, whose proof is deferred to Appendix C.1. The crucial step lies in the creation of the label space $\mathcal{Y}=R\cup S$, where $R$ is an infinite set whose points are all distance 2 apart, and $S$ is an infinite set whose elements are indexed by the finite subsets of $R$, e.g., as in $s_{R^{\prime}}\in S$ for finite $R^{\prime}\subseteq R$. For all such $R^{\prime}$, define $s_{R^{\prime}}$ to be distance 1 from the elements of $R^{\prime}$, distance 2 from the other elements of $R$, and distance 1 from all other points in $S$. Then $\mathcal{Y}$ indeed forms a metric space, and it is straightforward to see that, for instance, the class of all functions from a one-element set to $\mathcal{Y}$ is more difficult to learn than its finite projections (equivalently, finite subsets).
@@
The proof of Theorem 3.9 is deferred to Appendix C.2, but let us briefly sketch the main idea. Fix $n\in\mathbb{N}$, $S\in\mathcal{X}^{n}$, and set $\epsilon=\xi(n)$. Consider again the collection of functions $R=\mathcal{H}|_{S}$ and variables $L=\bigcup_{S^{\prime}\subseteq S,|S^{\prime}|=n-1}\mathcal{H}|_{S^{\prime}}$, as described in the proof of Theorem 3.6. By the premise of the theorem, for any finite subset $R^{\prime}\subseteq R$, there exists an assignment of variables $L\to\mathcal{Y}$ which maintains all functions in $R^{\prime}$ below $\epsilon$. Each such assignment induces an apportionment of error to each function in $R^{\prime}$, i.e., a vector of length $n$ with positive entries summing to $\epsilon$. For $r\in R^{\prime}$ depending upon variables $\ell_{1},\ldots,\ell_{n}$, this apportionment tracks the contribution of each $\ell_{i}$ to the output of $r$. The central technical step of the proof is to demonstrate that one can assign apportionments to each node $r\in R$ such that any finite subset of the apportionments can be satisfied by an assignment of variables $L\to\mathcal{Y}$. Then let $\ell=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ be a variable. We assign $\ell$ to the value $\hat{y}$ such that the function $(y_{1},\ldots,y_{i},\hat{y},y_{i+1},\ldots,y_{n})$ has minimal budget apportioned to $\ell$, among all such $\hat{y}$. From an invocation of the triangle inequality, this learner at most doubles the output of any $r\in R$.
@@
M. Hall’s theorem states that an infinite bipartite graph $G$ is $R$-matchable if and only if it is finitely $R$-matchable, provided that all nodes in $R$ have finite degree. Before proving M. Hall’s theorem by way of Theorem 3.3, we establish an intermediate lemma.
@@
The proof of Lemma 3.12 is deferred to Appendix C.3, but its intuition is fairly simple: by P. Hall’s theorem, $G$ is finitely $R$-matchable precisely when Hall’s condition holds, i.e., $|N(R^{\prime})|\geq|R^{\prime}|$ for all finite $R^{\prime}\subseteq R$ (Hall, 1935). Thus any $\ell\in L$ which is not incident to a Hall blocking set can be removed from $G$ while preserving Hall’s condition and finite $R$-matchability. Proceeding in this way, nodes can be removed until each remaining $\ell\in L$ is contained in a Hall blocking set $R^{\prime}_{\ell}$. At this point, $\ell$’s incident edges can be safely restricted to those which are incident with $R^{\prime}_{\ell}$, a finite set.
@@ 3.3 Agnostic Learning
Furthermore, one can use nearly identical reasoning as in the proofs of Theorems 3.6 and 3.7 to see that agnostic transductive learning is described by a system of variables $L$ and functions $R$. In particular, set $R=\mathcal{Y}^{n}$ and let $L\subseteq(\mathcal{Y}\cup\{?\})^{n}$ contain all sequences with exactly one ``?"``?"``?"` ` ? ". Then a function $r=(y_{1},\ldots,y_{n})$ depends upon the variables $\{\ell_{i}\}_{i\in[n]}$, where $\ell_{i}=(y_{1},\ldots,y_{i-1},?,y_{i+1},\ldots,y_{n})$ and $r(\ell_{1},\ldots,\ell_{n})=\frac{1}{n}\sum_{i\in[n]}d_{\mathcal{Y}}(y_{i},\ell_{i})-\inf_{h\in\mathcal{H}}\frac{1}{n}\sum_{i\in[n]}\ell(h(x_{i}),y_{i})$.
@@ 3.4 PAC Learning
Let $\mathcal{X}$ be a domain, $\mathcal{Y}$ a label set, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Suppose that the loss function $\ell$ is bounded and satisfies either of the following conditions: